OpenStudy (anonymous):

1. While simplifying some math work, Peter wrote on his paper that x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x3 + x3 + x3 + x3?

OpenStudy (nurali):

x^3⋅x^3⋅x^3⋅x^3=x^3+3+3+3=x^12 Now let's see the second part of the question. x^3+x^3+x^3+x^3=4x^3 In general, 4x^3≠x^12